If x is a standard normal variate find the following probabilities by using normal probability i) p(0.87<x<1.28)

ii) p(-0.34<x<0.62)

iii) p(x>0 85)

iv)p(x>-0.65)


1
Expert's answer
2021-07-12T16:05:34-0400

i)

"P(0.87<x<1.28)=P(x<1.28)-P(x\leq 0.87)"

"\approx0.8997274-0.8078498\approx0.091878"

ii)

"P(-0.34<x<0.62)=P(x<0.62)-P(x\leq -0.34)"

"\approx0.7323711-0.3669283\approx0.365443"

iii)

"P(x>0.85)=1-P(x\leq 0.85)"

"\approx0.197663"

iv)

"P(x>-0.65)=1-P(x\leq -0.65)"

"\approx0.742154"